# Kernelized Decoded Quantum Interferometry

### v1

@@ II Kernelized DQI: model and guarantees
This section formalizes kernelized decoded quantum interferometry (k‑DQI). The core idea is to insert a unitary kernel $K$ before the Fourier/Hadamard interference so that the spectrum of a shaped objective $P\!\circ\!f$ becomes more concentrated on decoder‑friendly supports. We quantify this concentration by a noise‑weighted head‑spectrum mass and prove that it monotonically controls the achievable approximation ratio in the standard DQI reduction‑to‑decoding pipeline [JordanNature2025]. Throughout, we reference coherent decoder modules and circuit constructions only for context [Patamawisut2025]; our results are theory‑ and simulation‑first and make no experimental claims.
@@ II.1 Pipeline and notation
$\widehat{\Sigma}_{K}(\ell,\eta;d):=\sum_{s\in S_{d}(\bm{\alpha})}\eta_{K}(s)\,|\alpha_{s}|^{2}\in[0,1].$ (3)
@@ II.2 The Monotonic Improvement Theorem
Physical Intuition & Implication: Theorem II.1 transforms the algorithm design problem from an abstract search for ”quantum advantage” into a concrete optimization task: maximize the spectral energy $\Sigma_{K}$. Intuitively, the kernel $K$ acts as a ”spectral lens.” Just as an optical lens focuses light to overcome transmission loss, a well-design kernel concentrates the problem’s solution mass into the few low-frequency modes (the ”head”) that classical decoders can reliably recover, even after attenuation by channel noise. (See Appendix A for the full operator-norm derivation and noise contraction bounds.)
@@ II.3 Noise model and decoder property
On the decoding side we only require a mild monotone-threshold property that abstracts algebraic soft/list decoders on RS-like families and BP/density-evolution guarantees on LDPC-like families [GuruswamiSudan1999, KoetterVardy2003, RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]. Informally, we assume that when a decoder $\mathrm{Dec}$ is fed with measurement outcomes supported on a head set $\mathcal{S}_{d}(\alpha)$ under local noise weights $\eta_{K}$, its success probability in achieving a target approximation ratio $\rho\in(0,1]$ is a nondecreasing function of the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$. Concretely, there exists a nondecreasing response curve $\Phi$ such that
$\Pr\!\bigl[\mathrm{Dec}\ \text{returns a solution of value}\ \geq\rho\bigr]\;\geq\;\Phi\!\bigl(\Sigma_{K}(\ell,\eta;d),\,d,\,\text{ensemble parameters}\bigr),$
where the mapping from decoder success to objective value follows the standard DQI reduction [JordanNature2025, JordanArxiv2024]. For RS-like (OPI) instances, $\Phi$ can be made explicit using BM/KV/GS decoders [GuruswamiSudan1999, KoetterVardy2003], and for LDPC-like ensembles it is given by density-evolution or EXIT-style analyses [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]; the corresponding formulas and coding-theoretic justification are collected in Appendix A.
@@ III.2 Limitation: delocalized spectra
Informally, a shaped spectrum is $\delta$-isotropic if its mass is spread almost uniformly across all $2^{n}$ modes: for a typical random direction $u$ orthogonal to any fixed $d$-dimensional subspace, the overlap $|\langle u,\alpha\rangle|^{2}$ is at most $\delta/2^{n}$, and the total mass in the top-$d$ order statistics is likewise bounded by roughly $d\,\delta/2^{n}$.
@@ IV.1 OPI / RS-like structure over finite fields
$\sum_{s\in\mathcal{S}_{d_{\star}}(\alpha)}|\alpha_{s}|^{2}\;\geq\;1-\varepsilon_{r,p},\qquad\varepsilon_{r,p}=o_{p\to\infty}(1)\ \text{for fixed }r,$
@@
The PPC property formalizes the intuition that quadratic-phase kernels “flatten” low-degree algebraic phases in the OPI pipeline, leaving only a small tail. Theorem IV.1 shows that this spectral sharpening translates into a controllable head-mass boost and, via $F_{\mathrm{RS}}$, into RS-style decoding gains that are robust under local noise through the parameters $\underline{\eta}$ and $\Delta(w)$. Figure 3 makes this effect visible: tuning a single kernel parameter suffices to cross the soft-decoder threshold and thereby strictly improve upon unkernelized DQI.
@@ IV.2 Sparse Max-XORSAT / LDPC-like structure
Let $B\in\{0,1\}^{m\times n}$ be a sparse parity-check matrix with left/right degree distributions $(\lambda,\rho)$ and girth at least $\Omega(\log n)$ on typical instances. The DQI reduction maps the shaped state $g(x)=P(f(x))$ through $\mathcal{F}$ to amplitudes supported on rows of $B$ and local combinations thereof (see [JordanNature2025, JordanArxiv2024]). We consider block-local kernels that act as the identity outside disjoint blocks of size $b=O(1)$ and preserve factor-graph locality.
@@
Unlike the OPI setting, LDPC-like ensembles do not admit global diagonalization by simple kernels. However, block-local alignment suffices to add head mass in a controlled, extensive way while preserving sparsity assumptions. Theorem IV.2, together with the finite-length and cross-ensemble thresholds in Figures 4 and 5, shows that such block-local kernels manifest as rightward shifts of DE/EXIT thresholds proportional to $\Delta\Sigma_{\mathrm{loc}}$, which, via the noise-aware lower bound, translates into improved k-DQI guarantees.
@@ VI.1 Regime I: Explicit Algebraic Structure
For problems defined by known polynomial constraints, such as the Optimal Polynomial Interpolation (OPI) problem discussed in Section III, the kernel parameters are analytically derivable. Given a target polynomial $\mathcal{P}$ of degree $\ell$, the optimal spectral shaping is achieved by matching the quadratic phase profile of the kernel to the curvature of $\mathcal{P}$. specifically, for a chirp kernel $K(\gamma)=\exp(-i\gamma\hat{x}^{2})$, the optimal rate satisfies $\gamma^{*}\propto\text{coeff}(\mathcal{P})$, which can be computed in $O(1)$ classical time. In this regime, the “search” cost is zero, and k-DQI provides a deterministic compilation advantage.
@@ VI.2 Regime II: Variational k-DQI for Latent Structure
Figure 8(b) validates the search efficiency. Even when initialized at a random configuration with negligible spectral overlap (white circle), a standard gradient-based optimizer reliably converges to the global optimum (red star) in $O(1)$ iterations. The trajectory demonstrates that the landscape gradients are robust against noise ($\eta=0.1$), and the absence of barren plateaus—guaranteed by the Lipschitz bounds derived in Appendix C—makes finding the optimal kernel parameters computationally scalable.
@@
Crucially, our data indicates that $\Sigma_{K}$ is robust to parameter misspecification: achieving 90% of the peak advantage only requires precision $\delta\theta\sim O(1/\sqrt{n})$, as analyzed in Appendix C.
@@ VII Discussion and outlook
Equally important is what the present framework does not allow. The isotropy upper bound formalizes a limitation often seen informally in the DQI discourse: when the shaped spectrum is delocalized, no unitary kernel can manufacture head mass at scale. This aligns our positive results with recent non-advantage observations and places k-DQI firmly on the conservative side of the advantage debate. Kernelization strengthens DQI where structure is present and learnable; it does not conjure structure where none exists.
@@ A.2 Head–tail interference and decoder coupling
Finally, the DQI reduction maps decoder success to an approximation ratio lower bound $A(\ell,\eta,K;d)$ (see [JordanNature2025, JordanArxiv2024]); since $\Phi$ is nondecreasing in its first argument and $\Delta(w)$ does not depend on $K$, the right‑hand side is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. If two kernels give strictly ordered head masses above threshold, the inequality is strict by the monotonicity of $\Phi$, establishing the claim.
@@ B.2 Per-mode contraction and leakage
proof B.1 (Proof sketch)
The channel $\mathcal{Z}$ contracts off-diagonal entries of the density matrix by a factor $(1-\eta)$ on each qubit, which in turn reduces diagonal probabilities by at least $(1-\eta)^{q(s)}$ where $q(s)$ is the number of qubits on which mode $s$ has nontrivial support in the $\mathcal{F}$-basis. The diagonal loss channel $\mathcal{L}$ rescales probabilities by $\tau(s)$. When $\mathcal{M}=I$ the claim follows by taking $\eta_{K}(s)=(1-\eta)^{q(s)}\tau(s)$ and summing over $s$ in the head.
For a general banded unitary $\mathcal{M}$ of width $w$, probability in $\mathcal{S}_{d}(\alpha)$ can be redistributed only within a $w$-neighborhood of each index. Unitarity bounds the total mass that can leave the head by a function $\Delta(w)$ that vanishes as $w\to 0$; see [Bu2025Noise] for a detailed construction. Collecting the three contributions yields the stated inequality.
@@ B.4 Noise-aware lower bound
By Proposition B.1, the post-channel head probability is at least the effective head mass $\Sigma_{K}(\ell,\eta;d)-\Delta(w)-\mu^{2}\|g\|_{2}^{2}$. The decoder monotonicity property (assumption (A3) and Section II.3) converts any increase in this effective head mass into a nondecreasing improvement of the decoder success probability. The standard DQI reduction from decoding success to objective value (e.g. [JordanNature2025, JordanArxiv2024]) then yields the lower bound $A(\ell,\eta,K;d)$ with a response curve $F$ that depends on the decoder family and problem ensemble (BM/KV/GS for RS-like codes [GuruswamiSudan1999, KoetterVardy2003], density evolution / EXIT for LDPC-like ensembles [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]). Monotonicity of $F$ in its second argument gives the final statement.

### v2

@@ I Introduction
The fragility of standard DQI can be phrased as a precise design question. Given a shaped spectrum $\widehat{\psi}$ and a noisy channel acting after the interference layer, the native spectral geometry must align with the capabilities of a restricted decoder family. In practice this alignment is often destroyed by hardware imperfections or slight instance perturbations. Our goal in this work is to replace this incidental alignment by an explicit, tunable degree of freedom: we seek efficiently implementable unitary kernels $K$ such that, under a realistic noise model and for a given decoder family, the resulting noise–weighted head mass $\widetilde{H}_{k}$ is provably increased and the corresponding approximation–ratio guarantees are improved. In other words, rather than asking whether an instance “has structure,” we ask which kernels expose sufficient spectral structure to place the system above a known decoding threshold.
@@ II.1 Pipeline and notation
The conceptual workflow of the proposed Kernelized Decoded Quantum Interferometry (k-DQI) framework is depicted in Fig. 1. Unlike standard approaches that rely solely on the intrinsic structure of the problem, our protocol introduces an explicit spectral engineering layer. Usefully viewed as a three-stage pipeline, it proceeds by first shaping the objective function $f$ into a polynomial representation $P(f)$, followed by a quantum evolution where a unitary kernel $K$ actively reshapes the probability distribution before noise exposure. Finally, the noisy measurement outcomes are processed by a classical decoder to recover the candidate solution $x^{*}$. This modular design allows us to theoretically quantify and optimize the robustness of the algorithm via the kernel spectrum $\Sigma_{K}$.
@@
$\Sigma_{K}(\ell,\eta;d):=\sum_{s\in S_{d}(\bm{\alpha})}\eta_{K}(s)\,|\alpha_{s}|^{2}\in[0,1].$ (4)
@@ II.2 The Monotonic Improvement Theorem
Physical Intuition & Implication: Theorem II.1 transforms the algorithm design problem from an abstract search for ”quantum advantage” into a concrete optimization task: maximize the spectral energy $\Sigma_{K}$. Intuitively, the kernel $K$ acts as a ”spectral lens.” Just as an optical lens focuses light to overcome transmission loss, a well-designed kernel concentrates the problem’s solution mass into the few low-frequency modes (the ”head”) that classical decoders can reliably recover, even after attenuation by channel noise. (See Appendix A for the full operator-norm derivation and noise contraction bounds.)
ed Refer to caption
@@ II.3 Noise model and decoder property
On the decoding side we only require a mild monotone-threshold property that abstracts algebraic soft/list decoders on RS-like families and BP/density-evolution guarantees on LDPC-like families [GuruswamiSudan1999, KoetterVardy2003, RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]. We assume that when a decoder $\mathrm{Dec}$ is fed with measurement outcomes supported on a head set $\mathcal{S}_{d}(\alpha)$ under local noise weights $\eta_{K}$, its success probability in achieving a target approximation ratio $\rho\in(0,1]$ is a nondecreasing function of the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$. Concretely, there exists a nondecreasing response curve $\Phi$ such that
$\Pr\!\bigl[\mathrm{Dec}\ \text{returns a solution of value}\ \geq\rho\bigr]\;\geq\;\Phi\!\bigl(\Sigma_{K}(\ell,\eta;d),\,d,\,\text{ensemble parameters}\bigr),$ (7)
where the mapping from decoder success to objective value follows the standard DQI reduction [JordanNature2025]. For RS-like (OPI) instances, $\Phi$ can be made explicit using BM/KV/GS decoders [GuruswamiSudan1999, KoetterVardy2003], and for LDPC-like ensembles it is given by density-evolution or EXIT-style analyses [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]; the corresponding formulas and coding-theoretic justification are collected in Appendix A.
@@ III.2 Limitation: delocalized spectra
A shaped spectrum is $\delta$-isotropic if its mass is spread almost uniformly across all $2^{n}$ modes: for a typical random direction $u$ orthogonal to any fixed $d$-dimensional subspace, the overlap $|\langle u,\alpha\rangle|^{2}$ is at most $\delta/2^{n}$, and the total mass in the top-$d$ order statistics is likewise bounded by roughly $d\,\delta/2^{n}$.
@@ IV.1 OPI / RS-like structure over finite fields
$\sum_{s\in\mathcal{S}_{d_{\star}}(\alpha)}|\alpha_{s}|^{2}\;\geq\;1-\varepsilon_{r,p},\qquad\varepsilon_{r,p}=o_{p\to\infty}(1)\ \text{for fixed }r,$ (12)
@@
The PPC property formalizes the intuition that quadratic-phase kernels “flatten” low-degree algebraic phases in the OPI pipeline, leaving only a small tail. Eqs. 13 and 14 show that shows that this spectral sharpening translates into a controllable head-mass boost and, via $F_{\mathrm{RS}}$, into RS-style decoding gains that are robust under local noise through the parameters $\underline{\eta}$ and $\Delta(w)$. Figure 3 makes this effect visible: tuning a single kernel parameter suffices to cross the soft-decoder threshold and thereby strictly improve upon unkernelized DQI.
@@ IV.2 Sparse Max-XORSAT / LDPC-like structure
Let $B\in\{0,1\}^{m\times n}$ be a sparse parity-check matrix with left/right degree distributions $(\lambda,\rho)$ and girth at least $\Omega(\log n)$ on typical instances. The DQI reduction maps the shaped state $g(x)=P(f(x))$ through $\mathcal{F}$ to amplitudes supported on rows of $B$ and local combinations thereof (see [JordanNature2025]). We consider block-local kernels that act as the identity outside disjoint blocks of size $b=O(1)$ and preserve factor-graph locality.
@@
Unlike the OPI setting, LDPC-like ensembles do not admit global diagonalization by simple kernels. However, block-local alignment suffices to add head mass in a controlled, extensive way while preserving sparsity assumptions. The analysis above IV.2, together with the finite-length and cross-ensemble thresholds in Figures 4 and 5, shows that such block-local kernels manifest as rightward shifts of DE/EXIT thresholds proportional to $\Delta\Sigma_{\mathrm{loc}}$, which, via the noise-aware lower bound, translates into improved k-DQI guarantees.
@@ V.1 Resource model and module costs
Lemma V.1 (Synthesis of quadratic chirp kernels)
Let an nn–qubit index register encode $j\in\{0,\dots,2^{n}-1\}$ in binary, $j=\sum_{r=0}^{n-1}2^{r}j_{r}$ with $j_{r}\in\{0,1\}$. Consider the diagonal unitary
$|,K_{\theta}\;=\;\sum_{j=0}^{2^{n}-1}e^{i\theta j^{2}}\ket{j}\!\bra{j}$, (18)
which is a special case of the length–$N$ diagonal kernels in eq. 17, Then $K_{\theta}$ admits a decomposition into $O(n^{2})$ two–qubit controlled–phase gates and depth $O(n)$ in the Clifford+T model.
proof V.1
Expanding $j^{2}$ in terms of the binary digits yields
$j^{2}=\Bigl(\sum_{r=0}^{n-1}2^{r}j_{r}\Bigr)^{2}=\sum_{r=0}^{n-1}2^{2r}j_{r}\;+\;2\sum_{0\leq r<s\leq n-1}2^{r+s}j_{r}j_{s}.$ (19)
[…]
@@ VI.1 Regime I: Explicit Algebraic Structure
For problems defined by known polynomial constraints, such as the Optimal Polynomial Interpolation (OPI) problem discussed in Section IV, the kernel parameters are analytically derivable. Given a target polynomial $\mathcal{P}$ of degree $\ell$, the optimal spectral shaping is achieved by matching the quadratic phase profile of the kernel to the curvature of $\mathcal{P}$. specifically, for a chirp kernel $K(\gamma)=\exp(-i\gamma\hat{x}^{2})$, the optimal rate satisfies $\gamma^{*}\propto\text{coeff}(\mathcal{P})$, which can be computed in $O(1)$ classical time. In this regime, the “search” cost is zero, and k-DQI provides a deterministic compilation advantage.
@@ VI.2 Regime II: Variational k-DQI for Latent Structure
Figure 8(b) validates the search efficiency. Even when initialized at a random configuration with negligible spectral overlap (white circle), a standard gradient-based optimizer reliably converges to the global optimum (red star) in $O(1)$ iterations. The trajectory demonstrates that the landscape gradients are robust against noise ($\eta=0.1$), and the absence of barren plateaus—guaranteed by the Lipschitz bounds derived in Appendix E—makes finding the optimal kernel parameters computationally scalable.
@@
Crucially, our data indicates that $\Sigma_{K}$ is robust to parameter misspecification: achieving 90% of the peak advantage only requires precision $\delta\theta\sim O(1/\sqrt{n})$, as analyzed in Appendix E.
@@ VII Discussion and outlook
Equally important is what the present framework does not allow. The isotropy upper bound formalizes a limitation often seen in the DQI discourse: when the shaped spectrum is delocalized, no unitary kernel can manufacture head mass at scale. This aligns our positive results with recent non-advantage observations and places k-DQI firmly on the conservative side of the advantage debate. Kernelization strengthens DQI where structure is present and learnable; it does not conjure structure where none exists.
@@ A.2 Head–tail interference and decoder coupling
Finally, the DQI reduction maps decoder success to an approximation ratio lower bound $A(\ell,\eta,K;d)$ (see [JordanNature2025]); since $\Phi$ is nondecreasing in its first argument and $\Delta(w)$ does not depend on $K$, the right‑hand side is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. If two kernels give strictly ordered head masses above threshold, the inequality is strict by the monotonicity of $\Phi$, establishing the claim.
@@ B.2 Per-mode contraction and leakage
proof B.1
We treat the three components of the noise channel separately and then combine them.
(1) Local depolarizing and diagonal loss. Write the post–interference density matrix before noise as $|\rho=\ket{\widehat{\psi}}\!\bra{\widehat{\psi}}$. Local depolarizing of rate $p$ on each qubit contracts off–diagonal entries by a factor at most (1p)(1-p) per qubit. For a given spectral mode $i$, let $w_{Z}(i)$ denote the number of qubits on which the corresponding eigenvector has nontrivial support in the ZZ–basis. Standard estimates for product depolarizing channels imply that the diagonal entry $\rho_{ii}$ is reduced by at most a factor $(1-p)^{w_{Z}(i)}$; see, e.g., Ref. [Bu2025Noise] for a detailed derivation at the level of interference fringes. The diagonal loss channel with transmittances $\{t_{i}\}$ further rescales $\rho_{ii}$ by $t_{i}$, so that the combined effect of depolarizing and loss on mode $i$ is a multiplicative factor
$\lambda_{i}\;\geq\;(1-p)^{w_{Z}(i)}t_{i}.$ (30)
This defines the per–mode weights $\lambda_{i}$ appearing in the statement of the lemma.
(2) Banded unitary mixing. Let $U_{\mathrm{mix}}$ be a banded unitary of width $W$, meaning that each column of $U_{\mathrm{mix}}$ has support only on indices within a WW–neighborhood. Conjugating $\rho$ by $U_{\mathrm{mix}}$ can redistribute probability mass among indices, but unitarity and bandedness imply that mass initially in the head set $H_{k}$ can leak only into indices within a bounded neighborhood of $H_{k}$. In particular, the total probability that leaves $H_{k}$ due to $U_{\mathrm{mix}}$ is upper–bounded by a function $\delta_{\mathrm{mix}}(W)$ that vanishes as $W\to 0$ and remains $o(1)$ for fixed $W$ in the large–system limit; again see Ref. [Bu2025Noise] for an explicit construction.
(3) Combining the contributions. Let $P_{\mathrm{head}}$ denote the probability of measuring an index in $H_{k}$ after the full noise channel. From the discussion above we obtain
$P_{\mathrm{head}}\;\geq\;\sum_{i\in H_{k}}\lambda_{i}|\widehat{\psi}_{i}|^{2}\;-\;\delta_{\mathrm{mix}}(W),$ (31)
where the sum encodes the per–mode contraction from depolarizing and loss and the subtraction accounts for banded mixing. Absorbing $\delta_{\mathrm{mix}}(W)$ into the leakage term $\Delta$ in the lemma statement and using the normalization factor in the definition of the noise–weighted head mass $\widetilde{H}_{k}$ yields
$P_{\mathrm{head}}\;\geq\;\widetilde{H}_{k}(\bm{\lambda})-\Delta,$ (32)
with $\Delta\to 0$ as $W\to 0$. This is exactly the claimed inequality.
@@ B.4 Noise-aware lower bound
By Proposition B.1, the post-channel head probability is at least the effective head mass $\Sigma_{K}(\ell,\eta;d)-\Delta(w)-\mu^{2}\|g\|_{2}^{2}$. The decoder monotonicity property (assumption (A3) and Section II.3) converts any increase in this effective head mass into a nondecreasing improvement of the decoder success probability. The standard DQI reduction from decoding success to objective value (e.g. [JordanNature2025]) then yields the lower bound $A(\ell,\eta,K;d)$ with a response curve $F$ that depends on the decoder family and problem ensemble (BM/KV/GS for RS-like codes [GuruswamiSudan1999, KoetterVardy2003], density evolution / EXIT for LDPC-like ensembles [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]). Monotonicity of $F$ in its second argument gives the final statement.
@@ C.1 Formal PPC property
2. Explicit PPC for quadratic OPI with contiguous evaluation sets
The abstract PPC property of Definition C.1 is satisfied exactly (with zero tail) for a simple but representative family of OPI instances. This provides a concrete sanity check for the definition and shows that the constants can be made explicit in low-degree cases.
Lemma C.1 (Exact PPC for quadratic OPI over contiguous evaluation sets)
Let $p$ be an odd prime and consider an OPI instance over $\mathbb{F}_{p}$ with the canonical evaluation set $\{\alpha_{i}\}_{i=0}^{p-1}=\{0,1,\dots,p-1\}$ and an unknown quadratic polynomial
$h(x)=ax^{2}+bx+c,\qquad a\in\mathbb{F}_{p}^{\times},\;b,c\in\mathbb{F}_{p}.$ (37)
Suppose the shaping map in the DQI pipeline is chosen so that, up to normalization, the pre-kernel amplitudes are
$g_{x}\propto\exp\!\Bigl(\frac{2\pi i}{p}\,h(x)\Bigr),\qquad x\in\mathbb{F}_{p},$ (38)
and let $F_{p}$ denote the $p$-ary QFT. Define a quadratic chirp kernel
$|.K_{-a}:=\sum_{x=0}^{p-1}\exp\!\Bigl(-\frac{2\pi i}{p}\,ax^{2}\Bigr)\ket{x}\!\bra{x}$. (39)
Then the post-interference spectrum
$\alpha=F_{p}K_{-a}g/\|g\|_{2}$ (40)
is supported on a single mode: there exists a unique index $m^{\star}\in\{0,\dots,p-1\}$ such that $|\alpha_{m^{\star}}|^{2}=1$ and $\alpha_{m}=0$ for all $m\neq m^{\star}$. In particular, Definition C.1 holds with degree bound $r=2$, head size $d^{\star}=1$, and tail parameter $\varepsilon_{2,p}=0$ in this setting.
proof C.1
By construction of the shaping map and the kernel, the unnormalized post-kernel amplitudes are
$(K_{-a}g)_{x}\propto\exp\!\Bigl(\frac{2\pi i}{p}\,\bigl(ax^{2}+bx+c\bigr)\Bigr)\exp\!\Bigl(-\frac{2\pi i}{p}\,ax^{2}\Bigr)=\exp\!\Bigl(\frac{2\pi i}{p}\,(bx+c)\Bigr).$ (41)
Applying the $p$-ary QFT yields
$(F_{p}K_{-a}g)_{m}$ $\propto\sum_{x\in\mathbb{F}_{p}}\exp\!\Bigl(\frac{2\pi i}{p}\,(bx+c)\Bigr)\exp\!\Bigl(-\frac{2\pi i}{p}\,mx\Bigr)$ (42)
$=\exp\!\Bigl(\frac{2\pi i}{p}\,c\Bigr)\sum_{x\in\mathbb{F}_{p}}\exp\!\Bigl(\frac{2\pi i}{p}\,(b-m)x\Bigr).$ (43)
The inner sum is a finite geometric series. If $m\not\equiv b\pmod{p}$, then $\exp\!\bigl(2\pi i(b-m)/p\bigr)\neq 1$ and the sum vanishes. If $m\equiv b\pmod{p}$, every term in the sum equals $1$ and the sum evaluates to $p$. Thus there exists a unique index $m^{\star}\equiv b\pmod{p}$ for which $(F_{p}K_{-a}g)_{m^{\star}}\neq 0$, and $(F_{p}K_{-a}g)_{m}=0$ for all $m\neq m^{\star}$.
After normalization, the spectrum $\alpha=F_{p}K_{-a}g/\|g\|_{2}$ has $|\alpha_{m^{\star}}|^{2}=1$ and $\alpha_{m}=0$ for $m\neq m^{\star}$, so the entire spectral mass lies in the head of size $d^{\star}=1$ with no tail. This is exactly the PPC property of Definition C.1 with $r=2$, $d^{\star}=1$, and $\varepsilon_{2,p}=0$.
Corollary C.1 (Concrete PPC constants for quadratic OPI)
In the setting of Lemma C.1, the polynomial-phase concentration property holds with
$d^{\star}=1,\qquad\varepsilon_{2,p}=0.$ (44)
More generally, if the evaluation set consists of $m\leq p$ consecutive field elements and the shaping weights are not exactly uniform, one still obtains $\varepsilon_{2,p}=O(m^{-1})$ by a standard estimate on truncated geometric series.